OpenStudy (anonymous):

I need help on proving left side of the subtraction sigma notation to be equal to the right side. n n n Σ (ai - bi) = Σ ai - Σ bi i=m i=m i=m

OpenStudy (anonymous):

you can prove that by theorem

OpenStudy (anonymous):

if both ai and bi converge then you can apply the following rule which is what you've done ^_^

OpenStudy (anonymous):

. This what the teacher sent, You need to start with one side and end with the other side. It looks like you started with the right hand side and ended with the right hand side. I suggest starting with the left hand side and ending with the right hand side. The first few terms of the left hand side will be: (a1 - b1) + (a2 - b2) + (a3 + b3) +... Start with this and put all the a's together and put all the b's together. Work your algebra magic to get the right hand side of the equation.

OpenStudy (anonymous):

oh okay , you're on the right track so you'll have : (an - bn) where n can be any number :)

OpenStudy (anonymous):

in you case n = m = i :)

OpenStudy (anonymous):

it's a matter of sequence ^_^

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

(a1 + b1 ) - (a2 + b2 ).....(an + bn)

OpenStudy (anonymous):

is there anyway to get your email address so I can show you my scan work if it ok with you because it hard to type math work out

OpenStudy (anonymous):

sure, deviant.g@hotmail.com ^_^ I'll try to help out.

OpenStudy (anonymous):

Thanks I'll send you a scan copy of my work right now

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

your email address did no work here mine tariq_adediran@hotmail.com, so youcan just email me

OpenStudy (anonymous):

This is my work (a1+a2)+.....an-(b1+b2)+.......bin Which the teacher side that this is the right side and I should sart went the left side and end with the right side

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

sklee take the lead?

OpenStudy (anonymous):

You have \[\sum_{i=m}^{n} a_i - b_i\] on the left hand side, so it is (a_m - b_m) + (a_(m+1) - b_(m+1)) + ... + (a_(n-1) - b_(n-1)) + (a_n - b_n) = [a_m + a_(m+1) + ... + a_(n-1) + a_n] - [b_m + b_(m+1) + ... + b_(n-1) + b_n] = right hand side

OpenStudy (anonymous):

teriz, I'm sorry I'm in a hurry, sklee will explain it okay? :)

OpenStudy (anonymous):

That ok thanks anyway

OpenStudy (anonymous):

you're welcome , sklee lead please? ^_^

OpenStudy (anonymous):

sstarica, i will do my best

OpenStudy (anonymous):

awesome, thank you :) I very much appreciate it.

OpenStudy (anonymous):

you're welcome, have a good nite

OpenStudy (anonymous):

lol, it's morning, I'm in campus right now :) thank you

OpenStudy (anonymous):

oops..sorry, where are you at? I mean which country

OpenStudy (anonymous):

Hi Teriz, are you ok with that solution?

OpenStudy (anonymous):

Middle East :), alright gotta finish my assignment now ^_^

OpenStudy (anonymous):

ok, all the best sstarica :)

OpenStudy (anonymous):

thank you ^_^ likewise

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

I think so, did you just do the right side?

OpenStudy (anonymous):

oops I mean the left side?

OpenStudy (anonymous):

i started from the left to get the right hand side. I didnt type out the exact right hand side

OpenStudy (anonymous):

ya, i started from the left side

OpenStudy (anonymous):

ok, so will the right side be similar to the left except that when you you combine all the term

OpenStudy (anonymous):

ya, they are the same

OpenStudy (anonymous):

did you understand my teacher email to me because I still was kind of confuse about it?

OpenStudy (anonymous):

just use a simple illustration: m =1, n = 3 a_1 = 11, a_2 = 12, a_3 = 13; b_1= 1, b_2 = 2, b_3 = 3 The left hand side: (a_1 - b_1) + (a_2 - b_2) + (a_3 - b_3) = (11-1) + (12-2) + (13-3) = 10 + 10 + 10 = 30 The right hand side: [a_1 + a_2 + a_3] - [b_1 + b_2 + b_3] = [11 + 12 + 13] - [1 + 2 + 3] = 36 - 6 = 30

OpenStudy (anonymous):

which part that you don't understand?

OpenStudy (anonymous):

the left hand, because why is the minus inside for that one, but on the right hand sidethe minus is outside

OpenStudy (anonymous):

The left hand side means you find the difference between respective a and b terms before you sum them. On the right hand side, you sum the a and b terms before you find the difference

OpenStudy (anonymous):

So that was basically subtraction prove for the sum sigma notation you gave

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sklee is there anyone I can get your email address so I can show you my scan work tommrro

OpenStudy (anonymous):

sklin_04@hotmail.com

OpenStudy (anonymous):

Thanks for all help and by the way what college do you go to?

OpenStudy (anonymous):

i have graduated Teriz

OpenStudy (anonymous):

That good, so I'll email you my attached work sometime in the morning or afternoon tomrrow for you to check. Thanks again

OpenStudy (anonymous):

you're welcome,i will do my best

OpenStudy (anonymous):

ok goodnigt

OpenStudy (anonymous):

good nite